Logiweb codex of my factorial

Up Help my factorial, ∗!,

my factorial

[my factorial →^{pyk} “my factorial”]

∗!

[x! →^{val} If(x ≈ 0, 1, x · x − 1!)]

[x! →^{tex} “#1.!”]

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-16.UTC:17:53:35.048381 = MJD-53537.TAI:17:54:07.048381 = LGT-4625661247048381e-6